OpenStudy (anonymous):

Solve: 4|3w+1|-7=29 Will give medal and fan to shown work and answer.

OpenStudy (anonymous):

I am confused on what to do with the 4

OpenStudy (anonymous):

subtract 7 from both sides first, then divide 4 from both sides

OpenStudy (anonymous):

sorry add*

OpenStudy (anonymous):

\[ 4|3w+1|-7=29 \]is two equations. First would be: \[ 4[3w+1]-7=29 \]Second would be:\[ 4[-(3w+1)]-7=29 \]

OpenStudy (anonymous):

so after subtracting 7 from both sides i would do this: